OpenStudy (anonymous):

A flower shop received a total of 24 boxes of roses. Each box had red, yellow, and white roses. There were 16 roses in each box. The shop received a total of 125 red roses and 112 yellow roses. How many white roses did the shop receive?

OpenStudy (anonymous):

please help me u will get a medal

OpenStudy (lifeisadangerousgame):

Didn't I just do this question with you?

OpenStudy (anonymous):

no

OpenStudy (lifeisadangerousgame):

Yes...I did: http://openstudy.com/study#/updates/562a7742e4b09dcac88884c9

OpenStudy (anonymous):

no

OpenStudy (anonymous):

help plz

OpenStudy (anonymous):

do u know

OpenStudy (lifeisadangerousgame):

Sure. We already know that 24 boxes of roses were received, that there were a total of 16 roses inside each box, that there were 125 red roses altogether, and that there were 112 roses altogether. From just this information, we know how many roses there are altogether and eventually how many white roses the shop received. To get the amount of roses there are altogether, you multiply 24 boxes by 16 roses (inside each box. So: 24 * 16 = 384 So altogether there are 384 roses. Now, add the red and yellow roses together, so that you know how many roses you know you already have. 112 + 125 = 237 roses that are NOT white. Now subtract the total amount of roses that aren't white from the total amount of roses altogether: 384 - 237 = 147 white roses. Does that make sense?

OpenStudy (anonymous):

ty